OpenStudy (anonymous):

Prove that a triangle with sides x^2 - 1, 2x, and x^2 + 1 is a right triangle.

OpenStudy (anonymous):

What properties do you know for right triangles?

OpenStudy (anonymous):

You can use the pythagorean theroem in this problem.

OpenStudy (anonymous):

yep! a^2+b^2=c^2 Only problem here which one is a, b and c?

OpenStudy (anonymous):

X^2+1 is c because it's the largest.

OpenStudy (anonymous):

Maybe

OpenStudy (anonymous):

Isn't it the largest though out of the 3 expressions?

OpenStudy (anonymous):

what about 2x? Keep in mind that x can be any number

OpenStudy (anonymous):

(for this problem this part is not that important)

Parth (parthkohli):

No, no... from definition, the Pythagorean triplet is in form \(\rm 2x,x^2 - 1,x^2 + 1 \)

OpenStudy (anonymous):

so which is the hypotenuse?

Parth (parthkohli):

The hypotenuse is always \(\rm x^2 + 1\) in this case.

OpenStudy (anonymous):

@ParthKohli Thank you

Parth (parthkohli):

\[\rm (2x)^2 + (x^2 - 1)^2 = (x^2 +1)^2\]Test.

OpenStudy (anonymous):

OK that is right, but you do need to think about it! 2x can be equal to x^2+1 if x=1 (in this problem that makes no sense as x^2-1 would be 0)

Parth (parthkohli):

Not all Pythagorean triplets are in form \(\rm 2x,x^2 - 1 ,x^2 + 1 \), but all triplets in form \(\rm 2x , x^2 - 1, x^2 + 1\) are Pythagorean triplets.

Parth (parthkohli):

@Andras That does form a right angled triangle. Remember how \(\cos (0) = 1\)?

OpenStudy (raden):

just if x>=2

Parth (parthkohli):

It works even when \(\rm x = 1\).

Parth (parthkohli):

In general, always works in case \(\rm x \ge 1\).

OpenStudy (raden):

1^2-1 = 0

Parth (parthkohli):

Practically, it does form a triangle and that is how \(\cos(0) = 1\).

Parth (parthkohli):

By definition, it is not really a triangle, but is a valid figure.

OpenStudy (anonymous):

so this isn't a right triangle.

Parth (parthkohli):

Albeit, yes, you can stick with \(\rm x\ge 2\).

OpenStudy (anonymous):

@ParthKohli this isn't a right triangle right?

OpenStudy (anonymous):

because x =0 and u can't have a negative side.

Parth (parthkohli):

For now, it's not a right triangle, but theoretically it is (you'd learn this thing in trigonometry).

OpenStudy (raden):

maybe this link can help u (but in indonesian's language) http://androkurniawan.blogspot.com/2011/09/cara-sederhana-mendapatkan-bilangan.html

OpenStudy (shubhamsrg):

you may use cosine law..it'll automatically lead you the the fact that angle opposite to longest side is 90 degrees. longest side, you can easily see is x^2 + 1